# Using nonassociative algebras to classify skew polycyclic codes up to isometry and equivalence

### v1

@@ Abstract.
We propose new definitions of equivalence and isometry for skew polycyclic codes that will lead to tighter classifications than existing ones. This helps to reduce the number of previously known isometry and equivalence classes, and state precisely when these different notions coincide. In the process, we classify classes of skew $(f,\sigma,\delta)$-polycyclic codes with the same performance paramenters, to avoid duplicating already existing codes.
We exploit that the generator of a skew polycyclic code is in one-one correspondence with the generator of a principal left ideal in its ambient algebra. Algebra isomorphisms that preserve the Hamming distance (called isometries) map generators of principal left ideals to generators of principal left ideals and preserve length, dimension and Hamming distance of the codes. We allow the ambient algebras to be nonassociative, thus eliminating the need on restrictions on the length of the codes.
@@ Introduction
Let $S$ be a unital commutative associative ring and $f\in R=S[t;\sigma,\delta]$ a monic skew polynomial. When the context is clear, we will simply write skew polycyclic code instead of skew $(f,\sigma,\delta)$-polycyclic code, and skew constacyclic code instead of skew $(\sigma,c)$-constacyclic code (which is the when $f(t)=t^{m}-a$).
@@
2. We call two skew polycyclic codes isometric if there is a Hamming-weight preserving map between them that is canonically induced by a monomial isomorphism of degree $k$ between their ambient Petit algebras mapping $t$ to $\alpha t^{k}$ for some invertible $\alpha\in S^{\times}$ and restricting to some $\tau\in{\rm Aut}(S)$ over $S$, viewed as algebras over some subring of their center, where the Hamming weight of a polynomial $p(t)$ is defined as the number of nonzero coefficients of $p(t)$. Any such an isomorphism between Petit algebras, called an isometry, gives a correspondence between their ideals, hence between the associated codes, which preserves code rate and Hamming distance as well as length and dimension. This approach generalizes the previously used notions of isometry and equivalence between skew polycyclic codes that were defined using algebraic structures so far (which exclusively worked with isometries restricting to some $id\in{\rm Aut}(S)$), and is lead by our knowledge of isomorphisms between nonassociative Petit algebras [BrPum2019, Br2018, BrPumSt2018, BrPum2018, Pum2024].
@@
3. We compute all the classes of codes that are equivalent to a given one (Theorem 5.6). Our refined notion of equivalence yields tighter classifications, for some first examples for codes over finite fields see Section 5.2. (This was observed for the first time in [NevPum2025, Example 5.8].) We give a list of new and easy necessary and sufficient criteria for isometry and equivalence of skew polycyclic codes that will help distinguish non-equivalent codes with potentially different performance parameters.
@@
We will focus mostly on $(f,\sigma)$-polycyclic codes, especially skew constacyclic codes, as for these the isomorphisms and automorphisms of the corresponding ambient Petit algebras are well-known. By allowing the case that $\sigma=id$, we include the case that $R=S[t]$ where possible, thus recovering results from [CFLL2012, CDL2015, OuaHA25] as special cases. Overall, our equivalence relations will result in tighter partitions of non-equivalent classes of codes than the ones defined in [ALY2022, CON2025, OaNA25].
@@
(Skew) polycyclic codes over finite rings, e.g. over mixed alphabets, are among the most prominent linear codes currently studied. They can be used to construct quantum error-correcting codes, usually via Gray maps and the CSS construction, see for instance [BagPa2025, YSP23, BDUBY, VPSI23, PPI25]. We expect that the results of this paper will make the search for new SPC codes and correspondingly quantum error-correcting codes more efficient and more organized. Good parameters for quantum codes are still under investigation, and any results on isometries and equivalences for SPC codes can be used to optimize search algorithms, in turn facilitating the discovery of new record-breaking quantum codes. In particular, when considering the CSS construction for quantum error-correcting codes, one needs to concentrate on codes $C$ such that the dual code $C^{\perp}$ satisfies $C^{\perp}\subset C$ as explained in [DinhBagUpadhyayBandiTansuchat2021]. Thus an important next step in the classification will be to consider when isometries preserve some suitable notion of duality of codes.
@@
Our main results can be found in Section 5 onwards and employ Hamming weight preserving isomorphisms between Petit algebras treated in Section 4. We look at the case that $\delta=0$ and $\sigma$ an automorphism.
@@
we call two classes of skew $(f,\sigma)$ and $(h,\sigma)$-polycyclic codes of length $m$ isometric if there is an algebra isomorphism $G_{\tau,\alpha,k}$ between their ambient algebras $R/Rf$ and $R/Rh$ which restricts to some $\tau\in{\rm Aut}(S)$ and maps $t$ to $\alpha t^{k}$ for some integer $k$ with 0<k<m10<k<m-1 and some $\alpha\in S^{\times}$, and Chen isometric if $\tau=id$. When $k=1$, we call the classes equivalent, and Chen equivalent, when additionally $\tau=id$, and write $G_{\tau,\alpha}=G_{\tau,\alpha,k}$. The current literature only uses Chen isometry and Chen equivalence.
@@
(Theorem 5.2) Suppose $K$ is a field and $\mathrm{Aut}(K)$ is abelian. Let $n\geq m-1$ and assume that $f$ and $h$ do not generate two-sided ideals in $K[t;\sigma]$. Then two classes ${\bf C}_{f}$ and ${\bf C}_{h}$ of skew polycyclic codes of length $m$ over $S$ are equivalent if and only if they are isometric, and Chen equivalent if and only if they are Chen-isometric.
Note that it is known that if $S_{0}$ denotes the ring fixed by $\sigma$, and $a,b\in S\setminus S_{0}$, or $a,b\in S_{0}^{\times}$ and $n\nmid m$ then equivalence and isometry coincide for the classes of skew $(\sigma,a)$-constacyclic and skew $(\sigma,b)$-constacyclic codes of length $m$ [NevPu2025, Corollary 4.5].
@@
Conditions for when two skew polycyclic codes cannot be equivalent, are given in Section 5.3.
@@ 2.2. Skew $(f,\sigma,\delta)$-polycyclic codes
A left (resp., right) linear code of length $m$ over $S$ is a left (resp., right) submodule of the $S$-module $S^{m}$. From now on, let $f(t)=t^{m}-\sum_{i=0}^{m-1}a_{i}t^{i}\in S[t;\sigma,\delta]$ be a reducible monic polynomial of degree $m>1$.
@@ 3.2. Our approach
(3.1) $G\left(\sum_{i=0}^{m-1}a_{i}t^{i}\right)=\sum_{i=0}^{m-1}\tau(a_{i})(\alpha t^{k})^{i}.$
@@
To motivate these notions we stress that the classifications we will achieve are tighter when employing the notions of equivalence and isometry (Section 5.2).
@@
In [OaHA25, Definition 4.1], for $K=\mathbb{F}_{q}$, two skew polynomials $f_{1},f_{2}\in K[t;\sigma]$ are called Hammming equivalent (resp. rank equivalent, if there is a $K$-vector space isomorphism $\varphi:K[t;\sigma]/K[t;\sigma]f_{1}\to K[t;\sigma]/K[t;\sigma]f_{2}$ such that
@@
    $\varphi$ preserves the Hamming distance $d_{H}(g,h)=d_{H}(\varphi(g),\varphi(h)),$ resp. the rank distance $d_{R}(g,h)=d_{H}(\varphi(g),\varphi(h))$.
Here, the rank distance is defined over arbtrary subfields of $K$.
@@ 5.1. Equivalence, and when it coincides with isometry
Let $\alpha\in S^{\times}$ and assume that $\tau\in{\rm Aut}(S)$ commutes with $\sigma$. Then two classes ${\bf C}_{f}$ and ${\bf C}_{h}$ of skew polycyclic codes of length $m$ over $S$ are equivalent if and only if there exists $\tau\in{\rm Aut}(S)$ that commutes with $\sigma$ and $\alpha\in S^{\times}$ such that
@@
When we only consider skew constacyclic codes then we can say more.
@@
[NevPu2025, Corollary 4.5]
Then the classes of skew $(\sigma,a)$-constacyclic and skew $(\sigma,b)$-constacyclic codes of length $m$ over $S$ are equivalent if and only if there exists some $\tau\in{\rm Aut}(S)$ that commutes with $\sigma$ and some $\alpha\in S^{\times}$ such that
(5.1) $\tau(a)=N_{m}^{\sigma}(\alpha)b,$
and are are Chen equivalent if and only if $\tau=id$.
If $a,b\in S\setminus S_{0}$, or $a,b\in S_{0}^{\times}$ and $n\nmid m$ then equivalence and isometry coincide.
@@
5.2. Skew polycyclic codes over finite fields
We can now get some straightforward results for skew polycyclic codes over finite fields. More detailed results will be obtained in a future paper.
@@ 5.2. Skew polycyclic codes over finite fields
The next examples shows that using our refined notion of equivalence and isometry will give a tighter classification of constacyclic codes.
Example 5.11.
[NevPum2025, Example 5.8] Let $\gcd([m]_{s},p^{r}-1)=p^{r}-1$. Then $N_{m}^{\sigma}(\mathbb{F}_{p^{r}}^{\times})=\{1\}$ and so no two classes of skew $(\sigma,a)$-constacyclic and skew $(\sigma,b)$-constacyclic codes of length $m$ belonging to distinct $a,b\in\mathbb{F}_{p^{r}}^{\times}$ will be Chen equivalent. Thus, there are $|\mathbb{F}_{p^{r}}^{\times}|=p^{r}-1$ distinct classes of skew constacyclic codes under Chen equivalence.
Now there are at least $n$ choices for $\tau\in{\rm Gal}(\mathbb{F}_{p^{r}}/\mathbb{F}_{p})$ (corresponding to $\sigma^{\ell}$ for each $1\leq\ell<n$). This implies that the set $\{a^{p^{v}}\mid 0\leq v<r\}\subset\mathbb{F}_{q^{r}}$ is an equivalence class with $r$ elements. That is, all of the corresponding $(\sigma,a^{p^{v}})$-constacyclic codes of length $m$ are equivalent and so the set ${\bf C}_{t^{m}-a}]=\{{\bf C}_{t^{m}-a^{p^{v}}}\,|\,0\leq v<r\}$ has $r$ elements, while ${\bf C}_{t^{m}-a}]_{Chen}=\{{\bf C}_{t^{m}-a}\}$.
The following example which demonstrates how to apply the connection between skew polycyclic and skew constacyclic codes from Proposition 5.7 to a plethora of different scenarios and again clearly demonstrates that using our refined notion of equivalence gives a tighter classification of skew polycyclic codes.
Example 5.12.
Let $\gcd([m]_{s},p^{r}-1)=p^{r}-1$.
We construct the equivalence classes of some given $h(t)$ as in Theorem 5.6. We start with $h(t)=t^{m}-\sum_{i=0}^{m-1}b_{i}t^{i}$. For all $\alpha\in K^{\times}$, define
Then there exists a Chen equivalence $G_{id,\alpha}$ between $R/Rh_{id,\alpha}$ and $R/Rh$ and
By Proposition 5.7, employing Example 5.11, no two classes of skew $(f,\sigma)$-polycyclic and skew $(h,\sigma)$-polycyclic codes of length $m$ belonging to $f$ and $h$ with distinct constant coefficients $a_{0},b_{0}\in K^{\times}$ will be Chen-equivalent. Thus, there are at most $|K^{\times}|=p^{r}-1$ distinct classes of skew polycyclic codes of length $m$ under Chen equivalence.
Now take $\tau\in{\rm Gal}(\mathbb{F}_{p^{r}}/\mathbb{F}_{p})$ and set
then
${\bf C}_{h}]=\{{\bf C}_{h_{\tau,\alpha}}\,|\,\alpha\in\mathbb{F}_{p^{r}}^{\times},\tau\in{\rm Gal}(\mathbb{F}_{p^{r}}/\mathbb{F}_{p})\}.$
Note that the $h_{\tau,\alpha}(t)$ all have $\tau(b_{0})$ as constant coefficients, for some $\tau\in{\rm Gal}(\mathbb{F}_{p^{r}}/\mathbb{F}_{p})$, since $N_{m}^{\sigma}(K^{\times})=\{1\}$, and that there are at least $n$ choices for $\tau$.
This argument works more generally, too.
Example 5.13.
For $i\in\{0,\ldots,m-1\}$ set $m_{i}=m-i$. Suppose there is $i\in\{0,\ldots,m-1\}$ such that $\gcd([m_{i}]_{s},p^{r}-1)=p^{r}-1$. By Proposition 5.7, no two classes of skew $(f,\sigma)$-polycyclic and skew $(h,\sigma)$-polycyclic codes of length $m$ where $f$ and $h$ have distinct $a_{i_{0}},b_{i_{0}}\in\mathbb{F}_{p^{r}}^{\times}$ will be Chen-equivalent. Thus, there are again at most $|K^{\times}|=p^{r}-1$ distinct classes of skew polycyclic codes under $m$-Chen equivalence; one for each choice of $a_{i_{0}}\in\mathbb{F}_{p^{r}}^{\times}$. Note that every ${\bf C}_{h_{id,\alpha}}\in[{\bf C}_{h}]_{Chen}$ has $b_{i_{0}}$ as $i_{0}$-th coefficient. Every ${\bf C}_{h_{\tau,\alpha}}\in[{\bf C}_{h}]$ has $\tau(b_{i_{0}})$ as $i_{0}$th coefficient.
@@
As a consequence we immediately obtain bounds for skew polycyclic codes.

### v2

@@ Abstract.
Employing isomorphisms between their ambient algebras, we propose new definitions of equivalence and isometry for skew polycyclic codes that will lead to tighter classifications than existing ones. This reduces the number of previously known isometry and equivalence classes. In the process, we classify classes of skew $(f,\sigma,\delta)$-polycyclic codes with the same performance parameters, to avoid duplicating already existing codes, and state precisely when different notions of equivalence coincide.
The generator of a skew polycyclic code is in one-one correspondence with the generator of a principal left ideal in its ambient algebra. We allow the ambient algebras to be nonassociative, thus eliminating the need on restrictions on the length of the codes. Algebra isomorphisms that preserve the Hamming distance (called isometries) map generators of principal left ideals to generators of principal left ideals and preserve length, dimension and Hamming distance of the codes.
@@ Introduction
Let $S$ be a unital commutative associative ring and $f\in R=S[t;\sigma,\delta]$ a monic skew polynomial. When the context is clear, we will simply write skew polycyclic code instead of skew $(f,\sigma,\delta)$-polycyclic code, and skew constacyclic code instead of skew $(\sigma,c)$-constacyclic code (when $f(t)=t^{m}-a$). The most important case here will likely be the one where $S$ is a finite field or ring, however, the theory holds for any commutative unital ring and so we will develop it as such.
@@
2. We call two skew polycyclic codes isometric, if there is a Hamming-weight preserving map between them that is canonically induced by a monomial isomorphism of degree $k$ between their ambient Petit algebras mapping $t$ to $\alpha t^{k}$ for some invertible $\alpha\in S^{\times}$ and restricting to some $\tau\in{\rm Aut}(S)$ over $S$, viewed as algebras over some subring of their center, where the Hamming weight of a polynomial $p(t)$ is defined as the number of nonzero coefficients of $p(t)$. Any such an isomorphism between Petit algebras, called an isometry, gives a correspondence between their ideals, hence between the associated codes, which preserves code rate and Hamming distance as well as length and dimension. This approach generalizes the previously used notions of isometry and equivalence between skew polycyclic codes that were defined using algebraic structures so far (which exclusively worked with isometries restricting to some $id\in{\rm Aut}(S)$), and is lead by our knowledge of isomorphisms between nonassociative Petit algebras [BrPum2019, Br2018, BrPumSt2018, BrPum2018, Pum2024].
@@
3. We compute all the classes of codes that are equivalent to a given one (Theorem 5.7). Our refined notion of equivalence yields tighter classifications, for some first examples for codes over finite fields see Section 6. (This was observed for the first time in [NevPum2025, Example 5.8].) We give a list of new and easy necessary and sufficient criteria for isometry and equivalence of skew polycyclic codes that will help distinguish non-equivalent codes with potentially different performance parameters.
@@
We will focus mostly on $(f,\sigma)$-polycyclic codes, especially skew constacyclic codes, as for these the isomorphisms and automorphisms of the corresponding ambient Petit algebras are already rather well-known. By allowing the case that $\sigma=id$, we include the case that $R=S[t]$ where possible, thus recovering results from [CFLL2012, CDL2015, OuaHA25] as special cases. Overall, our equivalence relations will result in tighter partitions of non-equivalent classes of codes than the ones defined in [ALY2022, CON2025, OaNA25].
@@
(Skew) polycyclic codes over finite rings, e.g. over mixed alphabets, are among the most prominent linear codes currently studied. They can be used to construct quantum error-correcting codes, usually via Gray maps and the CSS construction, see for instance [BagPa2025, YSP23, BDUBY, VPSI23, PPI25]. We expect that the results of this paper will make the search for new SPC codes and correspondingly quantum error-correcting codes more efficient and more organized. Good parameters for quantum codes are still under investigation, and any results on isometries and equivalences for SPC codes can be used to optimize search algorithms, in turn facilitating the discovery of new record-breaking quantum codes. In particular, when considering the CSS construction for quantum error-correcting codes, one needs to concentrate on codes $C$ such that the dual code $C^{\perp}$ satisfies $C^{\perp}\subset C$ as explained in [DBUBT2021]. Thus an important next step in the classification will be to consider when isometries preserve some suitable notion of duality of codes.
@@
Our main results can be found in Section 5 onwards and employ Hamming weight preserving isomorphisms between Petit algebras treated in Section 4. We look at the case that $\delta=0$ and assume $\sigma$ is an automorphism.
@@
we call two classes of skew $(f,\sigma)$ and $(h,\sigma)$-polycyclic codes of length $m$ isometric, if there is an algebra isomorphism $G_{\tau,\alpha,k}$ between their ambient algebras $R/Rf$ and $R/Rh$ which restricts to some $\tau\in{\rm Aut}(S)$ and maps $t$ to $\alpha t^{k}$ for some integer $k$ with 0<k<m10<k<m-1 and some $\alpha\in S^{\times}$, and Chen isometric if $\tau=id$. When $k=1$, we call the classes equivalent, and Chen equivalent, when additionally $\tau=id$, and write $G_{\tau,\alpha}=G_{\tau,\alpha,k}$. The current literature only uses Chen isometry and Chen equivalence.
@@
(Theorem 5.2) Suppose $K$ is a field and $\mathrm{Aut}(K)$ is abelian. Let $n\geq m-1$ and assume that $f$ does not generate two-sided ideals in $K[t;\sigma]$. Then two classes ${\bf C}_{f}$ and ${\bf C}_{h}$ of skew polycyclic codes of length $m$ over $S$ are equivalent if and only if they are isometric, and Chen equivalent if and only if they are Chen isometric.
It is known that if $S_{0}$ denotes the ring fixed by $\sigma$, and $a,b\in S\setminus S_{0}$, or $a,b\in S_{0}^{\times}$ and $n\nmid m$ then equivalence and isometry coincide for the classes of skew $(\sigma,a)$-constacyclic and skew $(\sigma,b)$-constacyclic codes of length $m$ [NevPu2025, Corollary 4.5].
@@
Conditions for when two skew polycyclic codes cannot be equivalent, are given in Section 7.
@@ 2.2. Skew $(f,\sigma,\delta)$-polycyclic codes
A left (resp., right) linear code of length $m$ over $S$ is a left (resp., right) free submodule of the $S$-module $S^{m}$. From now on, let $f(t)=t^{m}-\sum_{i=0}^{m-1}a_{i}t^{i}\in S[t;\sigma,\delta]$ be a reducible monic polynomial of degree $m>1$.
@@ 3.2. Our approach
(3.1) $G\left(\sum_{i=0}^{m-1}c_{i}t^{i}\right)=\sum_{i=0}^{m-1}\tau(c_{i})(\alpha t^{k})^{i}.$
@@
To motivate these notions we stress that the classifications we will achieve are tighter when employing the notions of equivalence and isometry (Section 6).
@@
In [OaHA25, Definition 4.1], for $K=\mathbb{F}_{q}$, two skew polynomials $f_{1},f_{2}\in K[t;\sigma]$ are called Hamming equivalent (resp. rank equivalent), if there is a $K$-vector space isomorphism $\varphi:K[t;\sigma]/K[t;\sigma]f_{1}\to K[t;\sigma]/K[t;\sigma]f_{2}$ such that
@@
    $\varphi$ preserves the Hamming distance $d_{H}(g,h)=d_{H}(\varphi(g),\varphi(h)),$ resp. the rank distance $d_{R}(g,h)=d_{R}(\varphi(g),\varphi(h))$.
Here, the rank distance is defined over arbitrary subfields of $K$.
@@ 5.1. Equivalence, and when it coincides with isometry
(i) Two classes ${\bf C}_{f}$ and ${\bf C}_{h}$ of skew polycyclic codes of length $m$ over $S$ are equivalent if and only if there exists $\tau\in{\rm Aut}(S)$ that commutes with $\sigma$ and $\alpha\in S^{\times}$ such that
@@
(ii) The classes of skew $(\sigma,a)$-constacyclic and skew $(\sigma,b)$-constacyclic codes of length $m$ over $S$ are equivalent if and only if there exists some $\tau\in{\rm Aut}(S)$ that commutes with $\sigma$ and some $\alpha\in S^{\times}$ such that
(5.1) $\tau(a)=N_{m}^{\sigma}(\alpha)b,$
and are are Chen equivalent if and only if the above holds with $\tau=id$.
@@
Let $n\geq m-1$ and assume that $f$ does not generate two-sided ideals in $K[t;\sigma]$. Suppose $\mathrm{Aut}(K)$ is abelian. Then the notions of equivalence and isometry (resp., of Chen equivalence and Chen isometry) coincide for all skew $(f,\sigma)$-polycyclic codes over $S$ of length $m$.
This follows immediately from Theorem 4.2. When we only consider skew constacyclic codes then we can say more.
Corollary 5.3.
[NevPu2025, Corollary 4.5] Suppose $n\nmid m$, or that $a$ is not in $S_{0}$. Then the notions of equivalence and isometry (resp., of Chen equivalence and Chen isometry) coincide for skew $(\sigma,a)$-constacyclic codes over $S$ of length $m$.
@@
6. Skew polycyclic codes over finite fields
We can now get some straightforward results for skew constacyclic codes over finite fields. More detailed results will be obtained in a future paper.
@@ 6. Skew polycyclic codes over finite fields
The number of different Chen equivalence classes of skew constacyclic codes is $w=\gcd([m]_{s},p^{r}-1)$ [OaNA25]. Thus the number of polynomials that yield ambient algebras that are Chen equivalent to a fixed $t^{m}-a$ is $\frac{q-1}{\gcd([m]_{s},p^{r}-1)}$.
@@
The next examples show that using our refined notion of equivalence and isometry will give a tighter classification of constacyclic codes.
Example 6.2.
[NevPum2025, Example 5.8] Let $\gcd([m]_{s},p^{r}-1)=p^{r}-1$. Then $N_{m}^{\sigma}(\mathbb{F}_{p^{r}}^{\times})=\{1\}$ and so no two classes of skew $(\sigma,a)$-constacyclic and skew $(\sigma,b)$-constacyclic codes of length $m$ belonging to distinct $a,b\in\mathbb{F}_{p^{r}}^{\times}$ will be Chen equivalent. Thus, there are $|\mathbb{F}_{p^{r}}^{\times}|=p^{r}-1$ distinct classes of skew constacyclic codes under Chen equivalence. Each class is related to exactly one ambient algebra.
Moreover, if additionally $n\nmid m$ then Chen isometry and Chen equivalence coincide. Thus, there are also $p^{r}-1$ distinct classes of skew constacyclic codes under Chen isometry in that case.
When additionally $n|m$, then there are $\left(1-\frac{1}{[n]_{s}}\right)(p^{r}-1)$ different Chen isometry classes.
Now there are $r$ choices for $\tau\in{\rm Gal}(\mathbb{F}_{p^{r}}/\mathbb{F}_{p})$ ($n$ of which are corresponding to $\sigma^{\ell}$ for each $1\leq\ell<n$). This implies that the set $\{a^{p^{v}}\mid 0\leq v<r\}\subset\mathbb{F}_{q^{r}}$ is an equivalence class with exactly $r$ elements when $a\not\in\mathbb{F}_{p}$. That is, all of the corresponding $(\sigma,a^{p^{v}})$-constacyclic codes of length $m$ are equivalent and so the set ${\bf C}_{t^{m}-a}]=\{{\bf C}_{t^{m}-a^{p^{v}}}\,|\,0\leq v<r\}$ has $r$ elements when $a\not\in\mathbb{F}_{p}$, while ${\bf C}_{t^{m}-a}]_{Chen}=\{{\bf C}_{t^{m}-a}\}$. This yields a tighter classification of codes with the same Hamming distance, length and dimension.
Moreover, if $n\nmid m$ then isometry and equivalence coincide, so again calculating the isometry classes will not yield a tighter classification.
We count the number of equivalence classes in some special cases, to show how the notion of equivalence reduces the number of classes.
Example 6.3.
(communicated by M. Nevins)
Let $\gcd([m]_{s},q-1)=q-1$.
(i) Let $l$ be prime, $K=\mathbb{F}_{q}=\mathbb{F}_{p^{l^{2}}}$, $F=\mathbb{F}_{p^{l}}$ and $F_{0}=\mathbb{F}_{p}$. Then there are $p^{l^{2}}-p^{l}$ different $b\in K\setminus F$, and $l^{2}$ elements in each equivalence class belonging to one such $b$ (since $|{\rm Gal}(K/F_{0})|=l^{2}$), which yields
$\frac{p^{l^{2}}-p^{l}}{l^{2}}$
different equivalence classes. There are $p^{l}-p$ different $b\in F\setminus F_{0}$, $l$ elements in each equivalence class belonging to one such $b$ (since $|{\rm Gal}(F/F_{0})|=l$), which yields
$\frac{p^{l}-p}{l}$
different equivalence classes, and p1p-1 different $b\in F_{0}^{\times}$, yielding an additional p1p-1 different equivalence classes. We obtain a total of only
$\frac{p^{l^{2}}-p^{l}}{l^{2}}+\frac{p^{l}-p}{l}+p-1$
equivalence classes, compared with a total of
$p^{l^{2}}-1=(p^{l^{2}}-p^{l})+(p^{l}-p)+(p-1)$
equivalence classes with respect to Chen equivalence.
(ii) Let $K=\mathbb{F}_{p^{2}}$ and $F=\mathbb{F}_{p}$, i.e. assume $l=1$ in $(i)$. We know $\sigma(\xi)\not=\xi$ and $\mathbb{F}_{p^{2}}[t;\sigma]/\mathbb{F}_{p^{2}}[t;\sigma](t^{m}-a)\cong\mathbb{F}_{p^{2}}[t;\sigma]/\mathbb{F}_{p^{2}}[t;\sigma](t^{m}-\sigma(a))$ via $G_{id,1}$. However, $\mathbb{F}_{p^{2}}[t;\sigma]/\mathbb{F}_{p^{2}}[t;\sigma](t^{m}-a)$ and $\mathbb{F}_{p^{2}}[t;\sigma]/\mathbb{F}_{p^{2}}[t;\sigma](t^{m}-\sigma(a))$ are not Chen equivalent.
Example 6.4.
Let $K=\mathbb{F}_{3^{2}}$ and $F=\mathbb{F}_{3}$, so that $\sigma$ is the Frobenius automorphism and $n=2$. Using the results from the Magma computations performed for [OaNA25, Example 3] for $n\in\{1,\dots,10^{8}\}$ and Theorem 6.1 we count the different possible equivalence classes of skew constacyclic codes of length $m$ over $K=\mathbb{F}_{3^{2}}$. There are $\gcd([m]_{1},8)=\gcd(\frac{3^{m}-1}{2},8)$ different Chen equivalence classes.
(i) Let $m$ be odd. Then $\gcd([m]_{1},8)=1$ and there is exactly one Chen equivalence class and since $2\nmid m$ also exactly one Chen isometry class.
(ii) Let $m\equiv 2(\text{mod}\,4)$ then $\gcd([m]_{1},8)=4$ and there are 4 Chen equivalence and only $4-\frac{4}{[2]_{1}}=3$ Chen isometry classes (since $2|m$).
(iii) Let $m\equiv 0(\text{mod}\,4)$ then $\gcd([m]_{1},8)=8$ and we have 8 Chen equivalence classes and only $8-\frac{8}{[2]_{1}}=6$ Chen isometry classes (since $2|m$).
On the other hand, we have $\frac{6}{2}+2=5$ equivalence classes, and potentially even less isometry classes.
Example 6.5.
Let $K=\mathbb{F}_{2^{2}}$ and $F=\mathbb{F}_{2}$, so that $\sigma$ is the Frobenius automorphism and $n=2$. Let $m$ be even. Then $\gcd([m]_{1},3)=\gcd(2^{m}-1,3)=3$ and there are three Chen equivalence classes of skew constacyclic codes of length $m$ over $K=\mathbb{F}_{2^{2}}$, and $3(1-\frac{1}{3})=2$ Chen isometry classes (this fixes [OaNA25, Example 4] where it was claimed there was only one such class). On the other hand, there are two equivalence classes and potentially less isometry classes.
When $m$ is odd, we have $\gcd([m]_{1},3)=\gcd(2^{m}-1,3)=1$, so there exists only one Chen equivalence class and we are done.
The exact number of equivalence classes under Chen equivalence for codes of length $m$ over finite fields computed in [OaHA25, Theorem 4.20] is
$N=\frac{(q-1)^{m_{0}}}{lcm_{j\in\{i_{0},i_{1},\dots,i_{m_{0}}-1\}}\big(\frac{q-1}{gcd(|m-j|_{s},q-1)}\big)}$
where the $j$ run over all the $m_{0}$ different nonzero coefficients of $f$. (We note that this number is the same for codes over commutative rings, not just over finite fields.)
The previous arguments work more generally, too, and again clearly demonstrate that using our refined notion of equivalence gives a tighter classification of skew polycyclic codes.
Example 6.6.
Let $\gcd([m]_{s},p^{r}-1)=p^{r}-1$. We construct the equivalence classes of some given $h(t)$ as in Theorem 5.7. We start with $h(t)=t^{m}-\sum_{i=0}^{m-1}b_{i}t^{i}$. For all $\alpha\in K^{\times}$, define
Then there exists a Chen equivalence $G_{id,\alpha}$ between the ambient algebras $R/Rh_{id,\alpha}$ and $R/Rh$ and
By Proposition 5.8, employing Example 6.2, no two classes of skew $(f,\sigma)$-polycyclic and skew $(h,\sigma)$-polycyclic codes of length $m$ belonging to any two $f$ and $h$ with distinct constant coefficients $a_{0},b_{0}\in K^{\times}$ will be Chen equivalent since $N_{m}^{\sigma}(K^{\times})=\{1\}$.
Now take $\tau\in{\rm Gal}(\mathbb{F}_{p^{r}}/\mathbb{F}_{p})$ and set
then
${\bf C}_{h}]=\{{\bf C}_{h_{\tau,\alpha}}\,|\,\alpha\in\mathbb{F}_{p^{r}}^{\times},\tau\in{\rm Gal}(\mathbb{F}_{p^{r}}/\mathbb{F}_{p})\}$
and
${\bf C}_{h}]_{Chen}\subset[{\bf C}_{h}]$
Note that now the $h_{\tau,\alpha}(t)$ in the equivalence class of $h$ all have $\tau(b_{0})$ as constant coefficients, for some $\tau\in{\rm Gal}(\mathbb{F}_{p^{r}}/\mathbb{F}_{p})$. There are $r$ choices for $\tau$. So when $b_{0}\in\mathbb{F}_{q}\setminus\mathbb{F}_{p}$ we know that there are $r$ different $\tau(b_{0})$ and thus ${\bf C}_{h}]$ must be larger. This means there will be less equivalence classes.
Moreover, any two $f$ and $h$ whose non-zero constant coefficients $a_{0}$ and $b_{0}$ where the sets $\{a_{0},\tau(a_{0}),\dots,\tau^{r}(a_{0})\}$ and $\{b_{0},\tau(b_{0}),\dots,\tau^{r}(b_{0})\}$ have no elements in common, will not yield equivalent classes of codes.
We can apply the connection between skew polycyclic and skew constacyclic codes from Proposition 5.8 to a plethora of different scenarios.
Example 6.7.
For $i\in\{0,\ldots,m-1\}$ set $m_{i}=m-i$. Suppose there is $i\in\{0,\ldots,m-1\}$ such that $\gcd([m_{i}]_{s},p^{r}-1)=p^{r}-1$. By Proposition 5.8, no two classes of skew $(f,\sigma)$-polycyclic and skew $(h,\sigma)$-polycyclic codes of length $m$ where $f$ and $h$ have distinct $a_{i_{0}},b_{i_{0}}\in\mathbb{F}_{p^{r}}^{\times}$ will be Chen-equivalent. Thus, there are $|K^{\times}|=p^{r}-1$ distinct classes of skew polycyclic codes under $m$-Chen equivalence; one for each choice of $a_{i_{0}}\in\mathbb{F}_{p^{r}}^{\times}$. Note that every ${\bf C}_{h_{id,\alpha}}\in[{\bf C}_{h}]_{Chen}$ has $b_{i_{0}}$ as $i_{0}$-th coefficient. Every ${\bf C}_{h_{\tau,\alpha}}\in[{\bf C}_{h}]$ has $\tau(b_{i_{0}})$ as $i_{0}$th coefficient.
As a consequence of Theorem 5.2, the notions of Chen equivalence and Chen isometry coincide for all skew $(f,\sigma)$-polycyclic codes of length $m$.
Let $n\geq m-1$ and assume that $f$ does not generate a two-sided ideal in $\mathbb{F}_{q}[t;\sigma]$. Then the notions of Chen equivalence and Chen isometry coincide for all skew $(f,\sigma)$-polycyclic codes over $\mathbb{F}_{q}$ of length $m$.